# Zeroth-Order Optimization Finds Flat Minima

### v1

@@ 1 Introduction
[…]
@@ 1.1 Related Works
Sharpness-Aware Minimization. Prior studies of flat minima have mostly centered on first-order methods. Algorithms designed to find flat minima have achieved strong empirical success, including Entropy-SGD [15], stochastic weight averaging [42, 45], and sharpness-aware minimization (SAM) [33]. Similar methods to SAM were proposed in [92, 102], and their efficiency and performance were further enhanced in e.g., [49, 62, 26, 104, 5, 50]. Also inspired by the insights in Eq. (2), Zhang et al. [95] proposed a gradient-based method that effectively minimizes the smoothed loss. A recent work [78] introduced a method that replaces the SAM perturbation with zeroth-order estimation. In addition to the trace of Hessian used in this work, other notions of sharpness have also been studied in the literature. One such example is the largest eigenvalue of the Hessian matrix, which has been shown to be implicitly penalized by (S)GD with large learning rates [20, 7, 87, 67, 23, 2] and SAM [89, 11]. Li et al. [51] proved that SAM implicitly promotes balanced solutions on scale-invariant problems. A sharpness measure based on the largest gradient norm in a local neighborhood was proposed in [99]. Recently, Ahn et al. [3] provided a formal definition of flat minima and studied the convergence complexity of finding them. A local concept of flat minima was used, defining them as local minima that are also stationary points of the trace of Hessian evaluated at limit points under gradient flow. Two gradient-based algorithms were proposed with convergence guarantees to flat local minima under the assumptions that the loss function is four times continuously differentiable, satisfies the local PL condition, and has a twice Lipschitz limit map under gradient flow. In this work, we adopt a global notion of flat minima and assume convexity of the function to show that zeroth-order optimization with the two-point estimator converges to flat global minima.
@@ 3.1 Convergence Analysis
Summing up from $t=0$ to $t=T-1$ and averaging by $T$ give $\mathbb{E}[f(\bar{x}_{T})-\min_{x\in\mathbb{R}^{d}}f(x)]\leq\mathcal{O}(d/T)$ with a small enough $\lambda$, where $\bar{x}_{T}$ is the average of iterates [74].
@@
Our convergence guarantees require that $f(x)$ satisfies first-, second-, and third-order smoothness assumptions. In Appendix C, we explain how the first- and third-order smoothness assumptions can be relaxed, at the cost of slower convergence rates; see Table 1 for a summary.
@@ 4 Experiments
These two models can be viewed as linear probing [48] on a two-layer neural network with a frozen first layer and an identity activation. In the SVMs case, the second layer uses an identity activation and is trained with the squared hinge loss, while logistic regression applies a sigmoid activation and minimizes the cross-entropy loss. Both SVMs and logistic regression have convex objective functions, and the trace of Hessian can be efficiently computed. Figure 2 shows the optimization trajectories of gradient descent and zeroth-order optimization. In all cases, gradient descent and zeroth-order optimization achieve comparable training loss and test accuracy; however, zeroth-order optimization consistently reduces the trace of Hessian and converges to flatter solutions. Detailed experimental setups and additional results can be found in Appendix D.2.
@@ Acknowledgements
L.Z. gratefully acknowledges funding by the Max Planck ETH Center for Learning Systems. B.L. is supported by Swiss National Science Foundation Project Funding No. 200021-207343. This work does not relate to the current position of K.T. at Amazon. S.O. is supported in part by the National Science Foundation under grant no. 2019844, 2112471, and 2229876 supported in part by funds provided by the National Science Foundation, by the Department of Homeland Security, and by IBM. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation or its federal agency and industry partners. M.M. is supported by the German Research Foundation. N.H. is supported by ETH research grant funded through ETH Zurich Foundations and Swiss National Science Foundation Project Funding No. 200021-207343.
@@ Appendix B Missing Proofs from Section 3
which results in $(N-1)!!$! terms in the summation.
@@ B.1 Supporting Lemmas
$\lvert f(x+\lambda u)-f(x-\lambda u)\rvert$ $\leq\lvert f(x+\lambda u)-\Phi_{x}(x+\lambda u)\rvert+\lvert\Phi_{x}(x-\lambda u)-f(x-\lambda u)\rvert$
@@ B.3 Proofs of Theorem 1 and Corollary 2
We choose $\eta=1/(8(d+6)L_{1})$ such that $21-4\eta(d+6)L_{1}=1/$ 2. Rearranging terms, we have
@@
[…]
@@
Summing up from $t=0$ to $t=T-1$ and dividing both sides by $T$, we have that
@@
=\,= $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[F(x_{t})-F(x_{F}^{*})]$
[…]
@@ Appendix C Alternative Proof of the Main Result
$\lvert f(x+\lambda u)-f(x-\lambda u)\rvert$ $\leq\lvert f(x+\lambda u)-\Phi_{x}(x+\lambda u)\rvert+\lvert\Phi_{x}(x-\lambda u)-f(x-\lambda u)\rvert$
@@
We also briefly discuss how the $L_{1}$-smoothness assumption on $f(x)$ can be relaxed to the function being $L_{0}$-Lipschitz, i.e., $\lvert f(x)-f(y)\rvert\leq L_{0}\lVert x-y\rVert,\forall x,y\in\mathbb{R}^{d}$. By Lemma E.3 in [27], when $f(x)$ is convex and $L_{0}$-Lipschitz, $f_{\lambda}(x)$ is convex and $(L_{0}/\lambda)$-smooth. This means all dependence on $L_{1}$ can be replaced by $L_{0}/\lambda$ in the analysis below. With both second-order and third-order smoothness assumptions, and setting $\eta=\mathcal{O}(\lambda/d)$, this gives $\mathbb{E}[\textnormal{Tr}(\nabla^{2}f(\bar{x}_{T}))-\min_{x\in\mathcal{X}^{*}}\textnormal{Tr}(\nabla^{2}f(x))]\leq\mathcal{O}(d/(\lambda^{3}T)+\lambda^{3}d^{4}+\lambda^{2}d^{2})$ and thus $T=\mathcal{O}(\max\{d^{5}/\epsilon^{2},d^{4}/\epsilon^{5/2}\})$. With only second-order smoothness assumption, this gives $\mathbb{E}[\textnormal{Tr}(\nabla^{2}f(\bar{x}_{T}))-\min_{x\in\mathcal{X}^{*}}\textnormal{Tr}(\nabla^{2}f(x))]\leq\mathcal{O}(d/(\lambda^{3}T)+\lambda^{3}d^{4}+\lambda d^{3/2})$ and thus $T=\mathcal{O}(d^{11/2}/\epsilon^{4})$. Table 1 summarizes the complexity under different set of assumptions on $f(x)$.
@@
We choose $\eta=1/(8(d+2)L_{1})$ such that $21-4\eta(d+2)L_{1}=1/$ 2. Rearranging terms, we have
@@
Summing up from $t=0$ to $t=T-1$ and dividing both sides by $T$, we have that
@@ D.3 Fine-Tuning Language Models on Text Classification Tasks
We follow experiment settings in Malladi et al. [69] and consider few-shot fine-tuning on RoBERTa-Large [^3] https://huggingface.co/FacebookAI/roberta-large (355M parameters) [61] with $K=32$ and $K=256$ examples per class on 3 text classification tasks: SST-2444https://huggingface.co/datasets/stanfordnlp/sst2 and SST-$5$https://nlp.stanford.edu/sentiment/ [82] for sentiment classification, and TREC [^6] https://huggingface.co/datasets/CogComp/trec [84] for topic classification. RoBERTa is available under the MIT License [^7] https://github.com/facebookresearch/fairseq/blob/main/LICENSE. The datasets are commonly-used benchmarks that are publicly available for research purposes. Our implementation is based on the codebase provided by Malladi et al. [69] and uses the same prompts. Their code is released under the MIT License [^8] https://github.com/princeton-nlp/MeZO/blob/main/LICENSE. The SST-2, SST-5, and TREC datasets contain 2, 5, and 6 classes, respectively. We consider $K=256$ only for the SST-2 dataset, and $K=32$ for all 3 datasets. The training set is constructed by sampling $K$ examples per class from the original dataset, and the test set is built by randomly selecting $1,000$ examples from the original test dataset.

### v2

@@ 1 Introduction
We use $\lVert\cdot\rVert$ for the Euclidean norm of vectors and $[m]$ for the set $\{1,2,\cdots,m\}$. The trace of a square matrix $J$ is denoted by $\textnormal{Tr}(J)$. A standard Gaussian random vector in $\mathbb{R}^{d}$ is written as $v\sim\mathcal{N}(0,\mathrm{I}_{d})$ and satisfies $\mathbb{E}[vv^{\top}]=\mathrm{I}_{d}$. A function $p:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is convex if $\alpha\,p(x)+(1-\alpha)\,p(y)\geq p(\alpha x+(1-\alpha)y)$, $\forall x,y\in\mathbb{R}^{d}$ and $\alpha\in(0,1)$. A function $q:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $r$th-order smooth with $L_{r}>0$ if it is $r$ times differentiable and $\lVert\nabla^{r}q(x)-\nabla^{r}q(y)\rVert\leq L_{r}\lVert x-y\rVert$, where $\lVert\nabla^{r}q(x)\rVert:=\max_{s\in\mathbb{R}^{d},\lVert s\rVert\leq 1}\lvert\nabla^{r}q(x)[s]^{r}\rvert$ and $\nabla^{r}q(x)[s]^{r}:=\sum_{i_{1},\cdots,i_{r}\in[d]}(\partial^{r}q(x)/\partial x_{i_{1}}\cdots\partial x_{i_{r}})s_{i_{1}}\cdots s_{i_{r}}$. We say for brevity that $q(x)$ is $L_{1}$-smooth if $r=1$.
@@ 1.1 Related Works
Sharpness-Aware Minimization. Prior studies of flat minima have mostly centered on first-order methods. Algorithms designed to find flat minima have achieved strong empirical success, including Entropy-SGD [15], stochastic weight averaging [42, 45], and sharpness-aware minimization (SAM) [33]. Similar methods to SAM were proposed in [92, 104], and their efficiency and performance were further enhanced in e.g., [49, 62, 26, 106, 5, 50]. Also inspired by the insights in Eq. (2), Zhang et al. [97] proposed a gradient-based method that effectively minimizes the smoothed loss. Several recent studies [95, 78, 94] proposed modifying zeroth-order optimization and combining it with principles from SAM to explicitly promote flat minima. In contrast, our work focuses on understanding the implicit regularization effects inherent within the standard zeroth-order optimization. In addition to the trace of Hessian used in this work, other notions of sharpness have also been studied in the literature. One such example is the largest eigenvalue of the Hessian matrix, which has been shown to be implicitly penalized by (S)GD with large learning rates [20, 7, 87, 67, 23, 2] and SAM [89, 11]. Li et al. [51] proved that SAM implicitly promotes balanced solutions on scale-invariant problems. A sharpness measure based on the largest gradient norm in a local neighborhood was proposed in [101]. Recently, Ahn et al. [3] provided a formal definition of flat minima and studied the convergence complexity of finding them. A local concept of flat minima was used, defining them as local minima that are also stationary points of the trace of Hessian evaluated at limit points under gradient flow. Two gradient-based algorithms were proposed with convergence guarantees to flat local minima under the assumptions that the loss function is four times continuously differentiable, satisfies the local PL condition, and has a twice Lipschitz limit map under gradient flow. In this work, we adopt a global notion of flat minima and assume convexity of the function to show that zeroth-order optimization with the two-point estimator converges to flat global minima.
@@ 3.1 Convergence Analysis
Summing up from $t=0$ to t=T1t=T-1 and averaging by $T$ give $\mathbb{E}[f(\bar{x}_{T})-\min_{x\in\mathbb{R}^{d}}f(x)]\leq\mathcal{O}(d/T)$ with a small enough $\lambda$, where $\bar{x}_{T}$ is the average of iterates [74].
@@
Our convergence guarantees require that $f(x)$ is three times continuously differentiable and satisfies first-, second-, and third-order smoothness assumptions. In Appendix C, we explain how the first- and third-order smoothness assumptions can be relaxed, at the cost of slower convergence rates; see Table 1 for a summary.
@@ 4 Experiments
We provide empirical results on binary classification tasks with convex losses and language model fine-tuning tasks. Our code is available at https://github.com/Liang137/FlatZero.
@@
These two models can be viewed as linear probing [48] on a two‐layer neural network with a frozen first layer and an identity activation. In the SVMs case, the second layer uses an identity activation and is trained with the squared hinge loss, while logistic regression applies a sigmoid activation and minimizes the cross‐entropy loss. Both SVMs and logistic regression have convex objective functions, and the trace of Hessian can be efficiently computed. Figure 2 shows the optimization trajectories of gradient descent and zeroth-order optimization. In all cases, gradient descent and zeroth-order optimization achieve comparable training loss and test accuracy; however, zeroth-order optimization consistently reduces the trace of Hessian and converges to flatter solutions. Detailed experimental setups and additional results can be found in Appendix D.2.
@@ Acknowledgements
L.Z. gratefully acknowledges funding by the Max Planck ETH Center for Learning Systems. B.L. is supported by Swiss National Science Foundation Project Funding No. 200021-207343. This work does not relate to the current position of K.T. at Amazon. S.O. is supported in part by the National Science Foundation under grant no. 2112471, 2229876, and 2505865 supported in part by funds provided by the National Science Foundation, by the Department of Homeland Security, and by IBM. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation or its federal agency and industry partners. M.M. is supported by the German Research Foundation. N.H. is supported by ETH research grant funded through ETH Zurich Foundations and Swiss National Science Foundation Project Funding No. 200021-207343.
@@ Appendix B Missing Proofs from Section 3
which results in (N1)!!(N-1)!! terms in the summation.
@@ B.1 Supporting Lemmas
[…]
@@ B.3 Proofs of Theorem 1 and Corollary 2
We choose $\eta=1/(8(d+6)L_{1})$ such that $1-4\eta(d+6)L_{1}=1/2$. Rearranging terms, we have
@@
$\leq\,$ $\frac{\mathbb{E}\lVert x_{t}-x_{F}^{*}\rVert^{2}-\mathbb{E}\lVert x_{t+1}-x_{F}^{*}\rVert^{2}}{\eta}+\frac{L_{3}}{6}\lambda^{4}(d+4)^{2}+\frac{L_{2}^{2}}{3}\eta\lambda^{4}(d+6)^{4}+\frac{L_{3}^{2}}{288}\eta\lambda^{6}(d+10)^{5}$
$\leq\,$ $8(d+6)L_{1}(\mathbb{E}\lVert x_{t}-x_{F}^{*}\rVert^{2}-\mathbb{E}\lVert x_{t+1}-x_{F}^{*}\rVert^{2})$
@@
Summing up from $t=0$ to t=T1t=T-1 and dividing both sides by $T$, we have that
@@
$=\,$ $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[F(x_{t})-F(x_{F}^{*})]$
$\leq\,$ $\frac{8(d+6)L_{1}\lVert x_{0}-x_{F}^{*}\rVert^{2}}{T}+\frac{L_{3}}{6}\lambda^{4}(d+4)^{2}+\frac{L_{2}^{2}}{24L_{1}}\lambda^{4}(d+6)^{3}+\frac{L_{3}^{2}}{1152L_{1}}\lambda^{6}(d+10)^{4}.$
@@ Appendix C Alternative Proof of the Main Result
[…]
@@
We also briefly discuss how the $L_{1}$-smoothness assumption on $f(x)$ can be relaxed to the function being differentiable and $L_{0}$-Lipschitz, i.e., $\lVert\nabla f(x)\rVert\leq L_{0},\forall x\in\mathbb{R}^{d}$. By Lemma E.3 in [27], when $f(x)$ is convex and $L_{0}$-Lipschitz, $f_{\lambda}(x)$ is convex and $(L_{0}/\lambda)$-smooth. This means all dependence on $L_{1}$ can be replaced by $L_{0}/\lambda$ in the analysis below. With both second-order and third-order smoothness assumptions, and setting $\eta=\mathcal{O}(\lambda/d)$, this gives $\mathbb{E}[\textnormal{Tr}(\nabla^{2}f(\bar{x}_{T}))-\min_{x\in\mathcal{X}^{*}}\textnormal{Tr}(\nabla^{2}f(x))]\leq\mathcal{O}(d/(\lambda^{3}T)+\lambda^{3}d^{4}+\lambda^{2}d^{2})$ and thus $T=\mathcal{O}(\max\{d^{5}/\epsilon^{2},d^{4}/\epsilon^{5/2}\})$. With only second-order smoothness assumption, this gives $\mathbb{E}[\textnormal{Tr}(\nabla^{2}f(\bar{x}_{T}))-\min_{x\in\mathcal{X}^{*}}\textnormal{Tr}(\nabla^{2}f(x))]\leq\mathcal{O}(d/(\lambda^{3}T)+\lambda^{3}d^{4}+\lambda d^{3/2})$ and thus $T=\mathcal{O}(d^{11/2}/\epsilon^{4})$. Table 1 summarizes the complexity under different set of assumptions on $f(x)$.
@@
We choose $\eta=1/(8(d+2)L_{1})$ such that $1-4\eta(d+2)L_{1}=1/2$. Rearranging terms, we have
@@
Summing up from $t=0$ to t=T1t=T-1 and dividing both sides by $T$, we have that
@@ D.3 Fine-Tuning Language Models on Text Classification Tasks
We follow experiment settings in Malladi et al. [69] and consider few-shot fine-tuning on RoBERTa-Large [^3] https://huggingface.co/FacebookAI/roberta-large (355M parameters) [61] with $K=32$ and $K=256$ examples per class on 3 text classification tasks: SST-2444https://huggingface.co/datasets/stanfordnlp/sst2 and SST-5555https://nlp.stanford.edu/sentiment/ [82] for sentiment classification, and TREC [^6] https://huggingface.co/datasets/CogComp/trec [84] for topic classification. RoBERTa is available under the MIT License [^7] https://github.com/facebookresearch/fairseq/blob/main/LICENSE. The datasets are commonly-used benchmarks that are publicly available for research purposes. Our implementation is based on the codebase provided by Malladi et al. [69] and uses the same prompts. Their code is released under the MIT License [^8] https://github.com/princeton-nlp/MeZO/blob/main/LICENSE. The SST-2, SST-5, and TREC datasets contain 2, 5, and 6 classes, respectively. We consider $K=256$ only for the SST-2 dataset, and $K=32$ for all 3 datasets. The training set is constructed by sampling $K$ examples per class from the original dataset, and the test set is built by randomly selecting $1,000$ examples from the original test dataset.